​Order: D5 font size increased by 1 font size increased by 1 ¼ NS 240 ml IV to infuse over 24 hours. Using a micro drop​ factor, the initial flow rate was correctly set at 10​ mcgtt/min. The​ client's IV infiltrated and it took 5 hours to restart. After 8 hours 200 ml of IV fluid remains to be infused. The infusion is behind schedule. Compute the new flow rate in​ mcgtt/min so that the infusion will finish on time. Note that flow rates may not be adjusted by more than​ 20% of the original rate. To what rate would you reset the drip​ rate, or would you contact the​ prescriber?
Answer #1

1 ml = 60 mcgtt

The initial flow rate = 10 mcgtt/min. for 240 ml solution

(240 ml IV to infuse over 24 hours)

The initial rate = 240ml/24hour = 10ml/hour

Mcgtt (micro drops) in minutes = 10x 60(drop factor)/60 (minutes) = 10 mcgtt/minute

Remaining IV fluid after 8 hours = 200 ml

Remaining time = 16 hours (24 – 8 hours)

Flow rate = 200 ml /16 hours = 12.5ml/hours

In mcgtt; 12.5x 60(drop factor)/60 (minutes) = 12.5 mcgtt/minute

The 20% of initial flow rate = 10 x 20/100 = 2 ml and new rate should be 12 mcgtt/minute

The new calculated rate is 25% higher than the initial flow rate, so the prescriber’s opinion can be obtained before administration of fluids.